R = 1 KR 20, 23.9 kа vre79 V. R=12KR ER & 26-8-2 5- 8.4 Кл Кл 1 R6=2.akh
Answer #1

Ik R6 82 kr {R3 26k 5=72V WW 2.2ku > From the circut, R3 )|R4 € R3/1R4 = 8.2K || 8. 2K = 8.2KX8-2K - 4.1kn 8. 2K+8.2k so R3 1

From the circuit, » Îs- VG = 72 = GMA 12K Is = 6MA) I= VR 72 = 9MA - (I = amal From the circuit, Ig= Iti5 = per 6mtam It =

Apply Kul for the circuit It Rit V5 t Dj R6 - Vy=0 VT = PF (Rit R6) +V5 = 15m (1K+2.2k) +72 = 15 (3-2) +72 =48+72 VT = 120v